OpenStudy (anonymous):

could somebody help me with this?? solve 5/(5-p)-p/(5-p)=-2 I just started on this lesson and am not sure how to work through it, thanks!

OpenStudy (anonymous):

I think you move both the (5-p) to the other side with the -2

OpenStudy (anonymous):

the p^2?

OpenStudy (anonymous):

For everyone's reference \[\frac{ 5 }{ 5-p } - \frac{ p }{ 5-p } = -2\]

OpenStudy (skullpatrol):

@domu thank you :)

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

oh yes I know that